### v4

# Gauge Symmetry as the Origin of Spin–Orbit Coupling of Light

Abstract
Dual symmetry is an intrinsic property of Maxwell’s equations, corresponding to a global U(1) symmetry in vacuum, with helicity as the associated conserved quantity. In this paper, we investigate light propagation in a helical fiber system using a field-theoretical approach and introduce an effective gauge field that emerges from the localization of dual symmetry. We show that the helical trajectory of light rays reveals this gauge field as a manifestation of spin–orbit coupling. Although orbit–orbit coupling also arises in such systems, the spin–orbit interaction possesses deeper physical significance, as it originates from the intrinsic dual symmetry embedded in Maxwell’s equations.
preprint: APS/123-QED
I Introduction
Symmetry is a cornerstone of modern physics. According to Noether’s theorem, every continuous symmetry in a system corresponds to a conserved quantity. For the electromagnetic field, a particularly important symmetry is duality symmetry—the invariance of Maxwell’s equations in vacuum under the exchange of electric and magnetic fields [27]. This dual symmetry becomes especially meaningful in the presence of magnetic charges, where it can be formulated in a fully covariant, relativistic framework using four-vectors [47, 24, 22, 38, 28, 3]. Within this context emerges the concept of “electromagnetic democracy” [4, 6]—a viewpoint in which electric and magnetic fields are treated on equal footing. In vacuum, this symmetry is embodied by the discrete transformation $({\bf E},{\bf B})\rightarrow({\bf B},-{\bf E})$, under which Maxwell’s equations remain unchanged. More generally, this duality can be expressed as a $U(1)$ rotation in the real $\left({\bf E},{\bf B}\right)$ plane:
$\left(\begin{array}[]{c}{\bf E}\\ {\bf B}\end{array}\right)\rightarrow\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{array}\right)\left(\begin{array}[]{c}{\bf E}\\ {\bf B}\end{array}\right)=R(\theta)\left(\begin{array}[]{c}{\bf E}\\ {\bf B}\end{array}\right),$ (1)
where $\theta$ is the duality rotation angle. The classical "electromagnetic democracy" is given at $\theta=\pi/2$. Following Eq.(1), the electromagnetic field exhibits a global $U(1)$ symmetry, ensuring that helicity is conserved, as dictated by Noether’s theorem [13]. Of course, a natural question arises: what are the implications if the aforementioned dual transformation is localized? We will show that in helical fibers, the dual symmetry can indeed be localized in an optical-path-dependent manner, thereby ensuring the conservation of helicity [21]. We promote the global dual transformation in Eq. (1) to a local one in terms of a phase transformation of complex fields:
${\bf D}\pm i{\bf B}\rightarrow\text{e}^{\mp i\theta({\bf x})}({\bf D}\pm i{\bf B}),$ (2)
where $\theta({\bf x})$ is a function of coordinate $\bf x$}. In the presence of a medium, the dual transformation should replace $\mathbf{E}$ with the electric displacement field $\mathbf{D}$ to maintain the source-free condition. Moreover, the corresponding rotation angle now depends on spatial coordinates, meaning that the original global $U(1)$ symmetry may become a localU(1)local\,U(1) symmetry.
We carry out our analysis in helicity space, defined by a transverse basis relative to a chosen reference direction—typically aligned with the incident light’s propagation direction. Without loss of generality, we take this to be the $z$-axis. The helicity eigenvectors satisfy: $({\bf e}_{z}\cdot{\bf s}){\bf e}_{\pm}=\pm{\bf e}_{\pm}$ where ${\bf s}=\{s_{i}\}^{jk}=-i\epsilon_{ijk}$ is the adjoint representation of spin-1 matrices. In this helicity basis, the transverse components of the electric displacement and magnetic field are expressed as $\mathcal{D}_{\perp}=(D_{+},D_{-})^{T}$ and $\mathcal{B}_{\perp}=(B_{+},B_{-})^{T}$. Furthermore, we introduce the 2-vector wavefunctions of the photon in the transverse plane as [20]
$\psi^{\pm}=\mathcal{D}_{\perp}\pm i\sigma_{3}\mathcal{B}_{\perp},$ (3)
In vacuum, $\psi^{\pm}$ represent the components of positive and negative energy states in the helicity basis, respectively[29]. A negative-energy solution can be interpreted as an antiparticle traveling backward in time along its worldlines. However, since the photon is its own antiparticle, an antiphoton can be interpreted as a photon mode propagating in the opposite direction. In our previous analysis, we found that in general media, the helical state undergoes nontrivial transformations between these two modes and evolves independently as its own degree of freedom [46].
By introducing a four-component wave function, $\psi=(\psi^{+}\ \psi^{-})^{\mathrm{T}}$, Maxwell’s equations can be recast into an optical Dirac-like form when interactions with a medium are taken into account [20]. Under the local dual transformation (2), the four-component wave function transforms as:
$aligned\psi\rightarrow e^{-i\gamma^{0}\hat{S}_{z}\theta(\mathbf{x})}\psi,$ (4)
where the helicity operator $\hat{S}_{z}$ is defined as
$aligned\hat{S}_{z}=\left(\begin{array}[]{cc}\sigma_{3}&\\ &\sigma_{3}\end{array}\right).$ (5)
II Dual gauge field
According to the principles of gauge theory [45], maintaining local symmetry requires the introduction of a corresponding gauge field. In general, the infinitesimal variation of the derivative term under a local transformation can be expressed as
$\delta(\partial_{i}\psi_{\sigma})=-ig\partial_{i}(\theta(x)\psi_{\sigma}),$ (6)
where $g$ is some coupling coefficient. To preserve gauge invariance and compensate for this additional term, a gauge field $A_{i}$ is introduced, thereby promoting the ordinary derivative to a covariant derivative,
$\partial_{i}\rightarrow D_{i},\quad D_{i}=\partial_{i}+igA_{i},\;\rightarrow\;\hat{K}_{i}=\hat{k}_{i}+gA_{i},$ (7)
where $\hat{K}_{i}$ represents the covariant derivative in the momentum operator form. In this context, the gauge potential $A_{i}$ is associated with the spatial gradient of the material parameter, typically taking the form $\partial_{i}\ln\epsilon^{-1}$.
The total Lagrangian can thus be written as
$\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{A},$ (8)
where $\mathcal{L}_{0}$ represents the free photon part associated with global dual symmetry and can be constructed from the corresponding optical Dirac equation [20], while $\mathcal{L}_{A}$ accounts for the interaction between the photon and the gauge field, arising from the localization of the symmetry via $\theta(x)$.
For a spin-degenerate symmetric medium, excluding the spatial derivative terms of constitutive tensors yields the zeroth order Lagrangian,
$\mathcal{L}_{0}=\overline{\psi}(i\gamma^{0}\partial_{0}-\hat{m}_{0}+\gamma^{0}\bm{\gamma}\cdot\hat{\bf p}^{0})\psi,$ (9)
where $\overline{\psi}=\psi^{\dagger}\gamma^{0}$ is the conjugated field, $\gamma^{\mu}$ are the gamma matrices in the Dirac representation and $\hat{\mathbf{k}}=-i\bm{\nabla}$ is the momentum operator. $\hat{m}_{0}$ and $\hat{\bf p}^{0}=\{{\bf p}_{\perp}^{0},p^{0}_{z}\}$ denote the effective mass and momentum operators,
$\hat{m}_{0}=Q_{0}\hat{k}_{z}-\Delta\hat{m}_{0},\quad\Delta\hat{m}_{0}=\mathbf{q}\cdot\hat{\mathbf{k}}_{\perp}-\displaystyle{\frac{q_{0}}{\hat{k}_{z}}}\hat{k}_{+}\hat{k}_{-}$, (10)
$\hat{p}^{0}_{\pm}=Q_{\pm 2}\hat{k}_{z}-2Q_{\pm 1}\hat{k}_{\pm}+\displaystyle{\frac{q_{0}}{k_{z}}}\hat{k}_{\pm}^{2},\quad p_{z}^{0}=0.$ (11)
In the helicity space, the transverse components are $\hat{\mathbf{k}}_{\perp}=(\hat{k}_{+},\hat{k}_{-})^{T}$ with $\hat{k}_{\pm}=(\hat{k}_{x}\mp i\hat{k}_{y})/\sqrt{2}$; the multipole moments are related to inverse permittivity/permeability tensors $\epsilon^{-1}=\mu^{-1}=\{\Pi_{ij}\}$ through a unitary transformation $\hat{U}$,
$aligned\hat{U}^{-1}\Pi\hat{U}=\left(\begin{array}[]{ccc}Q_{0}&Q_{+2}&Q_{+1}\\ Q_{-2}&Q_{0}&Q_{-1}\\ Q_{-1}&Q_{+1}&q_{0}\end{array}\right),$ (12)
where the decomposition leads to two monopole moments $Q_{0}=\frac{1}{2}(\Pi_{11}+{\Pi}_{22})$ and $q_{0}=\Pi_{33}$, dipole $Q_{\pm 1}=\frac{1}{\sqrt{2}}(\Pi_{13}\mp i\Pi_{23})$ that form a dipole vector ${\bf q}=\{Q_{+1},Q_{-1}\}$, and quadrupole $Q_{\pm 2}=\frac{1}{2}\left[\left(\Pi_{11}-\Pi_{22}\right)\mp i\Pi_{12}\right]$, corresponding to the scalar (spin-0), vector (spin-1), and tensor (spin-2) modes, respectively[20].
By replacing the original momentum operator $\hat{k}_{i}$ in the Lagrangian $\mathcal{L}_{0}$ with $\hat{K}_{i}$, we obtain the total Lagrangian $\mathcal{L}$ that includes the contribution from the gauge field,
$\mathcal{L}=\overline{\psi}(i\gamma^{0}\partial_{0}-\hat{M}_{0}+\gamma^{0}\bm{\gamma}\cdot\hat{\bf P}^{0})\psi,$ (13)
where
$\hat{M}_{0}=Q_{0}\hat{K}_{z}-\Delta\hat{M}_{0},\quad\Delta\hat{m}_{0}=\mathbf{q}\cdot\hat{\mathbf{K}}_{\perp}-\displaystyle{\frac{q_{0}}{\hat{K}_{z}}}\hat{K}_{+}\hat{K}_{-}$, (14)
$\hat{P}^{0}_{\pm}=Q_{\pm 2}\hat{K}_{z}-2Q_{\pm 1}\hat{K}_{\pm}+\displaystyle{\frac{q_{0}}{K_{z}}}\hat{K}_{\pm}^{2},\quad P_{z}^{0}=0.$ (15)
[…]
$\hat{k}_{i}\rightarrow\hat{k}_{i}-\gamma^{0}S_{z}\partial_{i}\theta.$ (16)
To preserve gauge invariance, the gauge field must transform in a way that compensates for this coordinate-dependent variation, namely $A_{i}\rightarrow A_{i}+\partial_{i}\theta$. By comparing this result with the covariant derivative defined in Eq. (7), the coupling coefficient $g$ can be identified with the matrix $\gamma^{0}\hat{S}_{z}$, which corresponds to the helicity operator in the helicity eigenfunction representation (see Sec.III).
If we retain only the leading term, the interaction between the gauge field and the photon field can be expressed as
$\mathcal{L}_{A}=\overline{\psi}\{-\gamma^{0}\hat{S}_{z}(Q_{0}A_{z}+\bm{q}\cdot\bm{A})+\gamma^{\pm}\hat{S_{z}}(Q_{\pm 2}A_{z}-2Q_{\pm 1}A_{\pm})\}\psi.$ (17)
This result implies that when the medium is decomposed into tensor modes according to the concept of the original excitation, interactions naturally arise between these modes, photons, and the dual gauge fields.
Alternatively, one can consider photon dynamics in an optical medium without introducing a gauge field. Thus, the Lagrangian becomes
$\mathcal{L}=\mathcal{L}_{0}+\mathcal{L}_{Q}$ (18)
where $\mathcal{L}_{Q}$ captures terms involving the first-order spatial derivatives[46],
$\mathcal{L}_{Q}=\overline{\psi}(-\hat{m}_{q}+\gamma^{0}\bm{\gamma}\cdot\hat{\bf p}^{q})\psi,$ (19)
with
$\hat{m}_{q}$ $=$ $\hat{k}_{z}Q_{0}+\frac{i}{2}{\bm{\nabla}}_{\perp}\cdot\mathbf{q}-\hat{L}^{z}_{q}\hat{S}_{z},$ (20)
$\hat{p}^{q}_{\pm}$ $=$ $\hat{k}_{z}Q_{\pm 2}-2\hat{k}_{\pm}Q_{\pm 1}+\frac{\hat{k}_{\pm}q_{0}}{k_{z}}\hat{k}_{\pm},\quad\hat{p}^{q}_{z}=0.$ (21)
The operator $\hat{L}^{z}_{0}$ exhibits the characteristic features of orbital angular momentum and is defined as
$\hat{L}_{q}^{z}=\hat{\bf L}_{q}\cdot{\bf e}_{z}=\frac{1}{2}({\bm{\nabla}}_{\perp}\times\mathbf{q})\cdot\bm{e}_{z}.$ (22)
More generally, to account for the curved trajectory of light, we reformulate the Maxwell equation in a curvilinear coordinate with metric $g=\{g_{ij}\}$. Given the fact that curved space is equivalent to a spin-degenerate medium with the constitutive tensors $\epsilon_{g}$ and $\mu_{g}$ obeying the simple equality to the spatial metric, i.e., $\epsilon_{g}^{-1}=\mu_{g}^{-1}=g$, we can have effective constitutive tensors simply by matrix multiplication, $\epsilon\rightarrow\epsilon_{g}\epsilon$, $\mu\rightarrow\mu_{g}\mu$ [44, 46].
When the curl of the dipole moment vanishes—i.e., when $\hat{L}_{q}^{z}=0$—a trivial local gauge field arises. In this case, the gauge potential takes the pure-gauge form $A_{i}=iU^{\dagger}\nabla_{i}U$, where the transformation factor take $U=e^{-i\gamma^{0}\hat{S}_{z}\theta(\mathbf{x})}$. This corresponds to a vanishing gauge field strength, $F_{ij}=0$. By identifying Eqs. (17) and (19), we find that the following condition must hold:
$\hat{k}_{i}Q_{j}=\gamma^{0}\hat{S}_{z}Q_{j}A_{i},$ (23)
where $Q_{j}$ collects all components of the medium tensor appearing in Eq. (12). Solving this relation yields
$Q_{j}=Q_{j}^{c}\,e^{i\gamma^{0}\hat{S}_{z}\int A_{i}\,dx^{i}},$ (24)
where $Q_{j}^{c}$ denotes a constant amplitude. At the local scale, the perturbations induced in the medium by incident light depend on its polarization. Distinct polarization states generate dipole responses with different orientations, implying that the polarization of light governs the dipole-induced modification of the medium, expressed as
$\triangle Q_{j}\approx\lambda A_{i}\triangle x^{i},$ (25)
where $\lambda$ represents the polarization (helicity) index of the light.
In the following chapter, we will examine how the spatial rotation of the dipole term $\bm{q}$ along the helical trajectories of photons induces nontrivial gauge fields, and show that these dipole-driven effects constitute the physical origin of the gauge fields associated with spin–orbit coupling.
III Gauge field representation of spin-orbit coupling
Since the helical eigenstate of light serves as an independent propagation quantity under spin-degenerate conditions, we can formulate the $U(1)$ dual symmetry by taking the helical eigenstate as the fundamental element. Let $\psi_{i}$ denote the $i$th-component of a four-vector field $\psi$. The constraint imposed by helicity conservation allows the wavefunction to be decomposed into two opposite helicity states $\psi_{\sigma}$ with the helicities $\lambda=\pm 1$: $\psi_{+1}=(\psi_{1},0,0,\psi_{4})^{T}$, and $\psi_{-1}=(0,\psi_{2},\psi_{3},0)^{T}$. In this helicity representation, the coupling term $\gamma^{0}\hat{S}_{z}$ can be interpreted directly in terms of the helicity eigenvalue $\lambda$. Under a local dual transformation, the helicity component $\psi_{\lambda}$ transforms as
$\psi_{\lambda}\rightarrow e^{i\lambda\theta(\mathbf{x})}\psi_{\lambda},\quad\lambda=\pm 1.$ (26)
Accordingly, the transformation in Eq. (4) reflects a helicity rotation symmetry, which manifests as a $U(1)$ subgroup of the spin group $SU(2)$, in the form $e^{i\bm{\sigma}\cdot\mathbf{n}}$. As the topological charge $l$ does not affect the fundamental dual symmetry, we will set aside its discussion for now and revisit the case of vortex light at a later stage.
[…]
For photonic systems, the arc length parameter $s$ can be regarded as a time parameter in natural units, and the Lagrangian takes the form
$\mathcal{L}=\overline{\psi}(i\gamma^{0}D_{0}-\hat{m}_{0}+\gamma^{0}\bm{\gamma}\cdot\hat{\bf p}^{0})\psi,$ (27)
The covariant derivative $D_{0}$ with respect to time can equivalently be expressed in terms of the path length derivative $\hat{D}_{s}$. In this representation, the only nonvanishing component of the gauge field appears along the propagation path, yielding
$D_{s}=\partial_{s}+i\gamma^{0}\hat{S}_{z}{\mathcal{A}}_{s}.$ (28)
Obviously, under the local transformation introduced in Eq. (4), incorporating the gauge field transforms,
${\mathcal{A}}_{s}\rightarrow{\mathcal{A}}_{s}+\partial_{s}\theta(s),$ (29)
the Lagrangian keeps invariant. Accordingly, the effective Lagrangian associated with the gauge field component $A_{s}$ in the helicity eigenfunction representation $\psi_{\lambda}$ is
$\mathcal{L}_{A}=-\sum_{\lambda}\lambda\psi^{\dagger}_{\lambda}{\mathcal{A}}_{s}\psi_{\lambda},$ (30)
which differs from the standard quantum electrodynamics (QED) interaction. Within our optical–Dirac framework, we formulate a QED-like theory for photons, in which the photon is described by a Dirac-like spinor field $\psi_{\lambda}$, and the $U(1)$ gauge field is induced by the material system—specifically, the spatial derivatives of the dipole components $\bm{q}$ —with the coupling strength determined by the helicity $\lambda$ of the light.
To describe the local geometry of the optical path, the Frenet coordinate system based on a moving orthonormal frame is used [40, 1, 46]. In this setting, the induced material tensor parameterized by the multipole moments reduces to the constant components, $Q_{0}=1$, $Q_{\pm 2}=0$, and the spatially varying components, $Q_{\pm 1}=\mp i\tau r_{\pm}$, $q_{0}=1+\tau^{2}r^{2}$, where $r_{\pm}=(x\mp iy)/\sqrt{2}$, $\tau$ is the torsion of the optical path. The dipole field $\mathbf{q}$ corresponds to a rotation vector in real space, ${\bf q}=\bm{\Omega}_{\tau}\times{\bf x}$ with $\bm{\Omega}_{\tau}=\tau{\bf e}_{z}$. Here, $\tau$ is a slowly varying spatial parameter and may be treated adiabatically as a constant, leading to $\hat{k}_{+}Q_{+1}=\hat{k}_{-}Q_{-1}=0$. Similarly, $\hat{k}_{\pm}q_{0}\approx 0$, and $\hat{\mathbf{k}}_{\perp}\cdot\mathbf{q}=0$, implying $\bf q$} is a divergence-free vector.
In addition to the longitudinal propagation term $\hat{k}_{z}$ in the effective mass, two correction terms arise from light–medium interaction: $\Delta\hat{m}_{0}$, associated with the homogeneous medium response, and $\hat{m}_{q}$, originating from the non-commutative nature of the momentum operator with spatially varying constitutive tensors. Neglecting the paraxial term, the first correction takes the form $\Delta\hat{m}_{0}=\mathbf{q}\cdot\hat{\mathbf{k}}_{\perp}=\hat{L}_{z}\tau$, while the second is given by $\hat{m}_{q}=S_{z}\tau$, obtained from the curl relation $i\hat{\mathbf{k}}_{\perp}\times\mathbf{q}=2\bm{\Omega}_{\tau}$.
Accordingly, the medium-induced Lagrangian simplifies to
$\mathcal{L}_{Q}=\overline{\psi}\hat{S}_{z}\tau\psi=\sum_{\lambda}\psi^{\dagger}_{\lambda}\lambda\tau\psi_{\lambda},$ (31)
which describes the interaction between light’s polarization and its extrinsic orbital degrees of freedom [46]. This term, known as the spin–orbit coupling [32, 12], originates from the first-order derivative properties of the medium itself.
In the following, we demonstrate that Eq.(31) and Eq.(30) are equivalent representations in the case of helical optical fiber systems. Consider an initial polarization vector of the displacement field $\mathbf{D}$ aligned along the $x$-axis, expressed as ${\bf e}_{x}={\bf e}_{+}+{\bf e}_{-}$. Due to the geometric phase effect, the components of the right and left circular polarization acquire phase factors of opposite sign, ${\bf e}_{\pm}\rightarrow e^{\mp i\gamma(s)}{\bf e}_{\pm}$. Accordingly, we introduce a new set of polarization vectors by a rotation, $({\bf e}^{\prime}_{x},{\bf e}^{\prime}_{y})^{T}=R(\gamma(s))({\bf e}_{x},{\bf e}_{y})^{T}$, where $\gamma(s)$ denotes the Berry phase, corresponding to the rotation angle of the polarization plane [14, 41]. These rotated vectors, ${\bf e}^{\prime}_{x}$ and ${\bf e}^{\prime}_{y}$, can be identified with the transformed electric and magnetic field components, $\mathbf{D}^{\prime}$ and $\mathbf{B}^{\prime}$, respectively. As a result, the electric field, magnetic field, and propagation direction form a complete orthonormal tetrad, defining a three-dimensional local reference frame. Here, the propagation direction corresponds to that of the photon. Therefore, the dual transformation defined in Eq. (2) is formally analogous to a rotation, where the rotation angle $\theta(s)$ is naturally identified with the geometric phase $\gamma(s)$. This relationship can be expressed as [31]
$\theta(s)=\gamma(s)=\int\Omega(s)ds.$ (32)
where $\Omega(s)$ represents the angular velocity along the propagation path. Notably, when light propagates along a helical trajectory, the angular velocity corresponds to the torsion $\tau$ of the path, $\Omega(s)=\tau$ for a uniformly helical waveguide. The redefined field vectors $\mathbf{D}^{\prime}$ and $\mathbf{B}^{\prime}$ undergo parallel transport, which is equivalent to a transformation from a non-inertial reference frame to an inertial one [15, 31, 8].The gauge field serves as the connection governing parallel transport in this formulation, which leads to the vanishing of the effective gauge field in the inertial frame:
${\mathcal{A}}_{s}+\partial_{s}\gamma(s)=0\quad\Rightarrow\quad{\mathcal{A}}_{s}=-\tau.$ (33)
[…]
Since photons can also carry intrinsic orbital angular momentum (IOAM), $\hat{\bf L}=\hat{\bf x}\times\hat{\bf k}$, we generalize the photon wave function to a twisted optical field, written as $\psi^{+}e^{il\varphi}$ and $\psi^{-}e^{-il\varphi}$, where $l$ is the topological charge and $\varphi$ is the azimuthal angle [2]. For such a twisted light beam, an additional first-order correction term arises from the coupling between IOAM and EOAM, $\sim\hat{\bf L}\cdot\hat{\bf L}_{q}$, which contributes to the Lagrangian $\mathcal{L}_{0}$ at the same order as Eq. (31):
$\mathcal{L}_{L}=\overline{\psi}\hat{L}_{z}\tau\psi=\sum_{\sigma}\psi^{\dagger}_{\lambda_{\sigma}}l\tau\psi_{\lambda_{\sigma}},$ (34)
where $\psi_{l+1}\equiv(\psi_{1}e^{il\varphi},0,0,\psi_{4}e^{-il\varphi})^{T}$ and $\psi_{l-1}=(0,\psi_{2}e^{il\varphi},\psi_{3}e^{-il\varphi},0)^{T}$ represent the helicity eigenstates modified by a vortex phase. We denote these states collectively as $\psi_{\lambda_{\sigma}}$. Here, the total helicity $\lambda_{\sigma}$ is defined as
$\lambda_{\sigma}=\mathbf{J}\cdot\frac{\mathbf{k}}{k},$ (35)
which, for a vortex beam with topological charge $l$, becomes $\lambda_{\sigma}=(l+\sigma)$, where $\sigma=\pm 1$ denotes the spin eigenvalue of the photon.
A corresponding IOAM–EOAM coupling term also arises in the Lagrangian $\mathcal{L}_{A}$. To relate this term to the gauge field, however, the original dual symmetry must be generalized:
$\psi_{\lambda_{\sigma}}\rightarrow e^{-i\lambda_{\sigma}\theta(s)}\psi_{\lambda_{\sigma}}.$ (36)
This transformation represents a local rotation that acts not only on the polarization basis but also on the wavefront structure, effectively shifting the azimuthal phase from $\varphi$ to $\varphi-\theta(s)$. For the Lagrangian in Eq. (27) to remain invariant under this local symmetry transformation, the medium response $\delta Q_{i}$ must compensate for the non-vanishing commutator between the OAM operator $\hat{L}_{z}$ and the Lagrangian. Physically, this variation describes the angular momentum transfer between the photon and the medium, potentially incorporating contributions from induced multipole moments. Consequently, reassigning a term from $\mathcal{L}_{0}$ to $\mathcal{L}_{A}$ requires the introduction of a compensating contribution that accounts for the medium’s variation.
The interaction between the gauge field and the photon wavefunction in momentum representation can be expressed as
$\mathcal{L}_{A}=-\sum_{\sigma}\lambda_{\sigma}\psi_{\lambda_{\sigma}}^{\dagger}\dot{\mathbf{k}}\cdot\mathbf{A}\psi_{\lambda_{\sigma}}.$ (37)
which is consistent with the definition of angular velocity in momentum space, $\Omega(s)=-\mathbf{A}\cdot\dot{\mathbf{k}}$ [8]. This formalism naturally extends to momentum space through the introduction of a covariant derivative,
$aligned D_{s}^{\sigma}=\partial_{s}+i\lambda_{\sigma}{\mathcal{A}}_{s}\rightarrow\frac{dk^{i}}{ds}D_{i}^{\sigma}=\frac{dk^{i}}{ds}(\partial_{i}+i(\lambda_{\sigma}{\mathcal{A}}_{i})),$ (38)
Accordingly, the covariant derivative in momentum space is defined as
$D_{i}^{\sigma}=\partial_{i}+i\lambda_{\sigma}{\mathcal{A}}_{i},$ (39)
and its commutator yields
$D_{i}^{\sigma},D_{j}^{\sigma}]=i\lambda_{\sigma}{\mathcal{F}}_{ij}.$ (40)
[…]
From the perspective of the classical analogy to quantum mechanics, Eq. (37) can be interpreted as the spin–orbit interaction term. It can also be viewed as a Coriolis-like effect arising in a non-inertial reference frame co-moving with the wave [9, 8]. The corresponding Euler–Lagrange equations for the classical variables $\mathbf{k}$ and $\mathbf{r}$ reproduce the conventional ray trajectory. The inclusion of spin–orbit coupling introduces a first-order correction to the group velocity, resulting in the optical Hall effect [46, 12]:
[…]
The Berry phase can also be understood as a gauge-invariant Wilson loop in momentum space, arising from the integral of the field strength along a closed loop in momentum space[34, 7]. The Wilson loop is given by
$U_{g}=\exp[i\lambda_{\sigma}\oint_{\Gamma}dk^{i}A_{i}(k)]=\exp[i\lambda_{\sigma}\oint_{\Sigma}dS^{ij}F_{ij}(k)],$ (42)
where $\Gamma$ is a closed path in momentum space, the $\Sigma$ is the surface that spans the closed loop $\Gamma$, and the $dS^{ij}$ represents an area element.
IV Conclusion
We investigate the localization of the dual symmetry of the electromagnetic field in helical systems using a field-theoretical framework, in which a novel gauge field emerges naturally as a consequence of the local symmetry. The spin–orbit interaction arising during beam propagation along helical trajectories is formally equivalent to the coupling between the helicity and this emergent gauge field. This correspondence provides a unified gauge-theoretic interpretation of the spin Hall effect of light, where such phenomena can be viewed as analogues of axion–electromagnetic coupling [43, 23, 36, 42]. For vortex beams, corrections from orbit–orbit interactions remain significant, yet these are fundamentally distinct from spin–orbit coupling, which originates from a deeper intrinsic symmetry. This distinction provides a fresh perspective for differentiating between the spin and IOAM of light.
A particularly important result is Eq. (30). In conventional electrodynamics, it is well known that charge conservation arises naturally from the U(1)U(1)gauge symmetry of the electronic system, and that promoting this symmetry to a local one introduces the electromagnetic gauge potential—enabling interactions between electrons via photon exchange. In a completely analogous manner, the dual symmetry considered here is also a $U(1)$ symmetry, whose associated conserved quantity is helicity. Localizing this dual symmetry introduces a dual gauge field, implying that photons interact with each other through the exchange of quanta associated with this field. Consequently, our framework may provide a fundamentally new theoretical basis for describing interactions between light and matter.
Although the present work focuses primarily on reinterpreting the spin–orbit interaction, it is also valuable to explore the connection between non-Abelian gauge fields and the medium when this dual symmetry is extended in an isospin-like manner. Furthermore, it is interesting to consider the physical consequences of breaking this dual symmetry, such as the lifting of spin degeneracy and the emergence of a $\gamma^{5}$ term.
The work was supported by the National Key R&\&D Program of China through Grant No. 2020YFC2201400, and the National Natural Science Foundation of China through Grant No. 12375084.